OpenStudy (fanduekisses):

Is the derivative of f(x)=x+ 1/x f'(x)=1-x^-2 ?

OpenStudy (anonymous):

yes, although if you want a useful f0rm you would write \[f'(x)=1-\frac{1}{x^2}\]

OpenStudy (fanduekisses):

ok thanks ^_^

OpenStudy (fanduekisses):

i'm trying to find the critical numbers

OpenStudy (solomonzelman):

wait, maybe it is\[\frac{x+1}{x}\]

OpenStudy (anonymous):

well there you go then a perfectly good example of what i just said now it is easy to find the critical numbers oh yeah, maybe it is

OpenStudy (solomonzelman):

but the space after the first x, indicate the correctness of your interpretation, satelline.

OpenStudy (fanduekisses):

x is undefined when the denominator equals zero rihgt?

OpenStudy (anonymous):

actually i am going to stick with my first answer \[f'(x)=\frac{x^2-1}{x^2}\]

OpenStudy (anonymous):

yes, but i would not really call that a critical number your original function is not defined at 0 so it would be a miracle of the derivative was

OpenStudy (jhannybean):

The denominator tells you where you have a vertical asymptote, I believe.

OpenStudy (jhannybean):

Or yeah, where the function is undefined, because there is a VA there.

OpenStudy (solomonzelman):

yes.

OpenStudy (jhannybean):

So solve: \(x^2-1=0\) This wll give you your critical points.

OpenStudy (fanduekisses):

\[\pm1\]

OpenStudy (solomonzelman):

yes.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

again, that is why you want to write the derivative of \(\frac{1}{x}\) as \(-\frac{1}{x^2}\) rather than \(x^{-2}\) which is correct but does not help your cause at all

OpenStudy (anonymous):

similarly, the derivative of \(\sqrt{x}\) is \(\frac{1}{2\sqrt{x}}\) not \(\frac{1}{2}x^{-\frac{1}{2}}\)